Consider a lamina that occupies the region D bounded by the parabola x = 1 - y2 and the coordinate axes in the first quadrant with density function p (x, y) = y. Find the center of mass.


1
Expert's answer
2021-05-07T14:32:39-0400

Given, the region D is bounded by the parabola x = 1 - y2 and the coordinate axes in the first quadrant with density function p (x, y) = y.

Moments,

"M_x=\int\int yp(x,y)dA\newline=\int_{x=0}^1 \int_{y=0}^1 yp(x,y)dA\newline=\int_{x=0}^1 \int_{y=0}^1 y^2dydx\newline=\int_{x=0}^1[ \frac{y^3}{3}]_{0}^1dydx\newline=\frac{1}{3}\int_{x=0}^1dx\newline=\frac{1}{3}[ x]_{0}^1\newline=\frac{1}{3}\newline"

"M_y=\int\int xp(x,y)dA\newline=\int_{x=0}^1 \int_{y=0}^1 xydydx\newline=\int_{x=0}^1x[ \frac{y^2}{2}]_{0}^1dydx\newline=\frac{1}{2}\int_{x=0}^1xdx\newline=\frac{1}{2}[ \frac{x^2}{2}]_{0}^1\newline=\frac{1}{4}"

And,

Mass,m is given by,

"m=\int\int p(x,y)dA\newline=\int_{x=0}^1 \int_{y=0}^1 ydydx\newline=\int_{x=0}^1[ \frac{y^2}{2}]_{0}^1dydx\newline=\frac{1}{2}\int_{x=0}^1dx\newline=\frac{1}{2}[ x]_{0}^1\newline=\frac{1}{2}"

Let (x,y) be the center of mass.

"y=\frac{M_y}{m}\newline=\frac{\frac{1}{4}}{\frac{1}{2}}\newline=\frac{1}{2}\newline
x=\frac{M_x}{m}\newline=\frac{\frac{1}{3}}{\frac{1}{2}}\newline=\frac{2}{3}"

Thus, center mass is "(\frac{2}{3},\frac{1}{2})."